OpenStudy (anonymous):

How do i prove 1. arcsin(-x)=-arcsin(x) 2. arsin (x) + arcos(x)= Pie/2

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

a triangle will do for the second one

OpenStudy (solomonzelman):

are you allowed to use the fact that \(\large\color{green}{ {\rm Sin}(-x)=-{\rm Sin}(x) }\) ?

OpenStudy (misty1212):

site is not working so well, is it?

OpenStudy (misty1212):



OpenStudy (misty1212):

arcsin(x) = A arccos(x) = B \[A+B=\frac{\pi}{2}\]

OpenStudy (misty1212):

@SolomonZelman sure hope so, then inverse of an any odd function is odd

OpenStudy (solomonzelman):

yes, the site is awful for me... I mean it is pretty bad on daily basis (before the update, but now it is even worse).

OpenStudy (solomonzelman):

Oh, I became green.... bye

OpenStudy (misty1212):

what does it mean to become green?

Nnesha (nnesha):

LOL solo